OpenStudy (anonymous):

how do you know if f(x)=x^1/3 is differentiable at x=0?

OpenStudy (bahrom7893):

take the derivative and plug in x=0

OpenStudy (anonymous):

i dont think it is differentiable

OpenStudy (bahrom7893):

okay here: f'(x) = (1/3)x^(-2/3) = 1 / (3x^(2/3)) plug in x = 0, u end up with a 0 in the denominator, therefore.. no it's not differentiable, him is right. and that's how u test it.

OpenStudy (anonymous):

okay so if it is undefined for that point then it is not differentiable?

OpenStudy (bahrom7893):

yes

OpenStudy (anonymous):

yes. if the derivative is undefined at any point, then the function is not differentiable at that point.

OpenStudy (anonymous):

awesome...thanks guys!!

OpenStudy (anonymous):

if you look at the picture you will see why. \[y=\sqrt[3]{x}\] has a nice corner at (0,0)

OpenStudy (anonymous):

called a cusp , however a zero in the denominator doesn't always suggest non differentiability.

OpenStudy (anonymous):

it is possible that the differernce quotient has different limits from the RHS and LHS

OpenStudy (anonymous):

true...but since the numerator is non-zero, it does :)

OpenStudy (anonymous):

cusp. good word that...